OpenStudy (anonymous):

If \[Z ^{1/3} = 2 ( \cos \frac{ \pi }{ 9 } + i \sin \frac{ \pi }{ 9 }) \] Then z = ?

OpenStudy (anonymous):

\[\left( z ^{1/2} \right)^{2}=2^{2}\left( \cos \frac{ \pi }{ 9 }+isin \frac{ \pi }{9 } \right)^{2}\]and\[z=4\left( \cos ^{2} \frac{ \pi }{ 9 }+2icos \frac{ \pi }{ 9 }\sin \frac{ \pi }{ 9 }+\sin ^{2} \frac{ \pi }{ 9 }\right)=4\left( \cos ^{2}\frac{ \pi }{ 9 }+\sin ^{2}\frac{ \pi }{ 9 } \right)+4i\left( 2 \cos \frac{ \pi }{ 9 }\sin \frac{ \pi }{ 9 }\right)\]

OpenStudy (anonymous):

\[(z^{1/3})^3=(2(\cos \pi/9+i \sin \pi/9))^3\]

OpenStudy (anonymous):

\[z=4\left( 1 \right)+4i \left( \sin \frac{ 2\pi }{ 9 } \right)=4\left( 1+i \sin \frac{ 2\pi }{ 9 } \right)\]like this ??

OpenStudy (phi):

perhaps it is easiest to see in polar form: cos(x)+i*sin(x)= e^(ix) and (cos(x)+i*sin(x)^n= e^(i*n*x)

OpenStudy (anonymous):

opps!!! the third term is wrong it must be \[\cos ^{2}\frac{ \pi }{ 9 }-\sin ^{2}\frac{ \pi }{ 9 }\]

OpenStudy (anonymous):

guys I get the following: \[8 ( \cos ^{3} \left(\begin{matrix}\pi \\ 9\end{matrix}\right) + i \sin ^{3}\left(\begin{matrix}\pi \\ 9\end{matrix}\right))\] Is this incorrect?

OpenStudy (anonymous):

I got \[z=4\left( \cos \frac{ 2\pi }{ 9 }+i \sin \frac{ 2\pi }{ 9 } \right)=4e ^{i \frac{ 2\pi }{ 9 }}\]for applied the polar form by @phi

OpenStudy (phi):

yes, it is incorrect use (cos(x)+i*sin(x)^n= e^(i*n*x) = cos(n*x)+i*sin(n*x) with n= 3 and do not forget 2^3 out front

OpenStudy (phi):

pin is almost right, except he used n=2 (bad eyes?)

OpenStudy (anonymous):

nah, I rearrange the term befor to get cos(2x)=cos^2(x)-sin^2(x) and sin(2x)=2cos(x)sin(x) and become that answer XD

OpenStudy (phi):

you want \[ (Z^{\frac{1}{3}})^3 \]

OpenStudy (anonymous):

If I change to polar form first I'll get (e^(i(pi)/9))^2 that's equal

OpenStudy (anonymous):

lol, I think that z^(1/2) sry guy

OpenStudy (anonymous):

Perhaps \[2^{3}(\cos \frac{ \pi }{ 27 }+ isin \frac{ \pi }{ 27 })\]?

OpenStudy (anonymous):

I am lost!!!!!!!!

OpenStudy (anonymous):

sorry about confuseing you

OpenStudy (phi):

\[ Z ^{1/3} = 2 ( \cos \frac{ \pi }{ 9 } + i \sin \frac{ \pi }{ 9 })\] you want Z, so raise both sides to the 3rd power \[ (Z ^{1/3})^3 = 2^3 ( \cos \frac{ \pi }{ 9 } + i \sin \frac{ \pi }{ 9 })^3\] you get \[ Z = 8 ( \cos \frac{ \pi }{ 9 } + i \sin \frac{ \pi }{ 9 })^3\] now use (cos(x)+i*sin(x))^n = cos(n*x)+i*sin(n*x)

OpenStudy (anonymous):

What do you have as the final answer pinguine! The same as mine above?

OpenStudy (anonymous):

\[Z=8\left( \cos \frac{ \pi }{ 9 } +isin \frac{ \pi }{ 9 }\right)^{3}=8\left( e ^{-\frac{ i \pi }{ 9 }} \right)^{3}=8e ^{-\frac{ \pi }{ 3 }}\]I got this, tochange into the sine cos form it's\[Z=8\left( \cos \frac{ \pi }{ 3 } +\sin \frac{ \pi }{ 3 }\right)\]

OpenStudy (anonymous):

**** the power of exponential is all positive sry guy

OpenStudy (anonymous):

look at ths Euler's formula http://fermatslasttheorem.blogspot.com/2006/02/eulers-formula.html

OpenStudy (anonymous):

Thus also = \[4\sqrt{3}i\]

OpenStudy (anonymous):

How about you @phi ??

OpenStudy (phi):

Chem, you have difficulty interpreting the expression... pin is getting closer, but the "i" does not disappear. Look at my previous post and read it carefully.

OpenStudy (anonymous):

Phi I get the answer as: \[Z = 8( \cos \frac{ \pi }{ 3 } + i \sin \frac{ \pi }{ 3 })\]

OpenStudy (anonymous):

yes I see, sorry sir. Pay attention of phi ^^

OpenStudy (anonymous):

Is this not correct?

OpenStudy (phi):

yes, and we know cos(pi/3) is 1/2 and sin(pi/3) is sqrt(3)/2 so you could put it in numerical form

OpenStudy (anonymous):

My final answer in the form of a + bi: \[4 + 4\sqrt{3}i\] Does everyone agree???????????

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Thank you @ pinguine!

OpenStudy (anonymous):

Sorry about confuseing you :P try to look at Euler's formula sir

OpenStudy (anonymous):

I will do so. Thank you for your excellent knowledge and help!

OpenStudy (anonymous):

Im not excellent XD